OpenStudy (anonymous):

3.79(x+4.45)=4.17(x-7.53) x=

OpenStudy (anonymous):

\[3.79\times4.45+4.17\times 7.53=4.17\times x - 3.79\times x\]\[3.79\times4.45+4.17\times 7.53=(4.17 - 3.79)\times x\]\[(3.79\times4.45+4.17\times 7.53) \div (4.17 - 3.79)=x\]\[x=125.0394737\]

OpenStudy (anonymous):

that was not it.